OpenStudy (gabylovesyou):

If h(x) = 2x + 3, find the value of x if h(x) = 9

jimthompson5910 (jim_thompson5910):

This time, you replace h(x) with 9 and solve for x h(x) = 2x + 3 9 = 2x + 3 Tell me what you get

OpenStudy (gabylovesyou):

x = 3

OpenStudy (gabylovesyou):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you are 100% correct

jimthompson5910 (jim_thompson5910):

9 = 2x + 3 9 = 2*3 + 3 9 = 6 + 3 9 = 9 which is true

OpenStudy (gabylovesyou):

thank you!! :)

jimthompson5910 (jim_thompson5910):

np